### v2

@@ 8.3 Robotic Navigation and Manipulation
Recently, considerable efforts have been directed toward integrating VLMs for robotic navigation and manipulation, which are crucial for search and rescue (SaR) operations. Effective robotic navigation and manipulation in challenging environments require a precise understanding of physical properties and geometric relationships. Geometric awareness includes assessing whether an object can fit inside another or identifying appropriate cover, enabling robots to make informed decisions and interact safely with their surroundings. Physical grounding is essential for search and rescue operations and hazard detection. [51] introduce PHYSOBJECTS, a dataset incorporating physical concepts like mass, fragility, deformability, and density. They fine-tune a VLM on this dataset and propose a robotic manipulation framework where the LLM queries the VLM, receives responses, and formulates an execution plan. VLM-GroNav [36] integrates a VLM-based reasoning module with physical grounding to evaluate terrain properties like slipperiness and deformability. The framework is tested on two robots: Ghost Vision 60  [^11] https://www.ghostrobotics.io/vision-60 and Clearpath Husky  [^12] https://clearpathrobotics.com/husky-a300-unmanned-ground-vehicle-robot/. [4] integrate VLMs with Behavior Trees (BTs) to detect, identify, and recover from failures in robotic tasks. The VLM analyzes visual input to recognize failures, suggest missing conditions and update the BT to handle similar failures in the future. VLM-Social-Nav [157] leverages VLMs to interpret social context and compute a navigation cost, enabling robots to make socially aware decisions in human environments.
@@ 9 Open Challenges and Future Directions
9.1 Accuracy Vs Efficiency Tradeoff
@@
9.2 Agentic AI and Distributed VLMs @Edge
@@
9.3 On-the-fly Fine-Tuning of VLMs
@@ 9.4 Multimodal VLM @Edge
9.5 Reasoning for Cohesive Environment
@@
9.6 Hallucination
An emerging challenge in optimizing VLMs for edge environments is hallucination, where models produce confident but factually incorrect outputs. The hallucinations in VLMs can stem from:
Object Hallucination: The VLM can hallucinate objects not present in the scene. These arise due to contexts in the scene or vague text prompts. For instance, an image with a dining table, plates, and glasses could lead the VLM models to believe that there is food and drinks on the table. The BLIP model outputs "Dining Table with food and drinks" when there is no food or drink. Another example would be an image with Dogs in a park and a prompt to explain the image, the model generates, "It is a Dog Park," which can be confused with Dogs in a regular Park. The difference is in identifying the leash in regular parks vs unleashed dogs in a Dog park. Such intricate details might have been missed during the training process.
Relation Hallucination: The VLM model could hallucinate relations between objects in a given scene. For instance, a man and a dog in a scene could be hallucinated for "Man playing fetch with Dog".
Temporal Hallucination: The VLM model could hallucinate the actions performed in a scene. For instance, a woman standing on a cliff with a harness could be hallucinated as "Woman bungee jumping from the cliff".
Although not directly evaluated for hallucination, [204] show that post-training quantization distorts activation distributions, crucial for semantic alignment. This increases the risk of hallucinated outputs in multimodal tasks. [90] show that knowledge distillation (KD) leads to hallucination due to higher exposure bias and overfitting to teacher’s labels. [11] show that LoRA leads to hallucinations due to static pre-trained knowledge. These effects are particularly concerning in safety-critical applications like medical imaging or autonomous driving, where factual grounding is nonnegotiable.
Some studies explore methods to reduce hallucination in vision-language models. SERPENT-VLM [76] introduces a self-supervised loss between the pooled image representation and the contextual representation of the generated text, which aligns vision and language features to reduce hallucination. This strategy can also be incorporated in compressed settings to maintain semantic and textual consistency. AdaVIB [8] mitigates hallucinations by using a variational information bottleneck that introduces stochastic noise to improve vision language alignment. Jiang et al. [69] suggest editing the model’s latent representation to reduce hallucination by 27% for LLaVa models. Some recent methods use decoding techniques [7] [44] [196] to mitigate the hallucination. As vision-language models are increasingly used in edge and distributed settings, future research should focus on developing compression-aware grounding objectives and uncertainty estimation techniques to mitigate hallucination. Models such as Grounded-SAM [148] help reduce object hallucinations as they require users to specify objects of interest, which may be more suitable for edge environments and applications that specifically look for known objects in the scene.
9.7 Security and Privacy

### v3

@@ 8.3 Robotic Navigation and Manipulation
Recently, considerable efforts have been directed toward integrating VLMs for robotic navigation and manipulation, which are crucial for search and rescue (SaR) operations. Effective robotic navigation and manipulation in challenging environments require a precise understanding of physical properties and geometric relationships. Geometric awareness includes assessing whether an object can fit inside another or identifying appropriate cover, enabling robots to make informed decisions and interact safely with their surroundings. Physical grounding is essential for search and rescue operations and hazard detection. [42] introduces PHYSOBJECTS, a dataset incorporating physical concepts like mass, fragility, deformability, and density. They fine-tune a VLM on this dataset and propose a robotic manipulation framework where the LLM queries the VLM, receives responses, and formulates an execution plan. VLM-GroNav [28] integrates a VLM-based reasoning module with physical grounding to evaluate terrain properties like slipperiness and deformability. The framework is tested on two robots: Ghost Vision 60  [^11] https://www.ghostrobotics.io/vision-60 and Clearpath Husky  [^12] https://clearpathrobotics.com/husky-a300-unmanned-ground-vehicle-robot/. [4] integrate VLMs with Behavior Trees (BTs) to detect, identify, and recover from failures in robotic tasks. The VLM analyzes visual input to recognize failures, suggest missing conditions and update the BT to handle similar failures in the future. VLM-Social-Nav [127] leverages VLMs to interpret social context and compute a navigation cost, enabling robots to make socially aware decisions in human environments.
@@ 9 Open Challenges and Future Directions
9.0.1 Accuracy Vs Efficiency Tradeoff
@@
9.0.2 Agentic AI and Distributed VLMs @Edge
@@
9.0.3 On-the-fly Fine-Tuning of VLMs
@@ 9.0.4 Multimodal VLM @Edge
9.0.5 Reasoning for Cohesive Environment
@@
9.0.6 Security and privacy